# Microphone Array Geometry Independent Multi-Talker Distant ASR: NTT System for the DASR Task of the CHiME-8 Challenge

### v1

@@ 5 Speech recognition backend
Figure 9: Our proposed ASR backend.
This section describes how we implemented our ASR backend “$\text{ASR}(\cdot)$” in Eq. (5). Figure 9 shows the schematic diagram of our ASR backend, which uses four different ASR models, performs language model (LM) for restoring, and combines their output with system combination. The details of the ASR models, the training data used for their training, the language model (LM) used for rescoring, and the system combination approach are described below.
@@ 5.1 Training data for ASR backend
We used 70 hours of CHiME-8 training data processed with GSS for the oracle segmentation, commonly used for building ASR models (ASR1-4). We did not use train_call and train_intv in Mixer 6 at this time because preliminary experiments showed minimal improvements with them. For the pre-training of the transducer-based ASR model with WavLM Large (ASR4), we also used LibriSpeech [55], and VoxCeleb1+2 [51] datasets in addition to the CHiME-8 described above. The development set was used for measuring token accuracy or training loss for early stopping.
@@ 5.2 ASR models
We employed four ASR models: two attentional encoder-decoder (AED) models [69] and two transducer models [70]. For building strong ASR systems, we adopt several pre-trained models: 1) two different sizes of Whisper-based AED models (ASR1/ASR2), 2) the official pre-trained transducer model based on the CHiME-8 DASR NeMo baseline system (ASR3), and 3) the self-supervised learning (SSL) model, i.e., WavLM [49], for the encoder of the transducer model (ASR4). The following describes the detailed model structure and training pipeline of each ASR model.
@@ 5.2.1 Whisper Large v3 (ASR1)
Whisper Large V3 model [71] is a strong ASR model trained on a large amount of data. It consists of 1540M parameters, i.e., 32-layer Transformer encoder and decoder, each with 8 attention heads and a hidden dimension of 1280. It uses the byte-level byte-pair encoding (BPE) [72] tokenizer from GPT-2 [73] and has a vocabulary size of 51,864 tokens. We finetuned all parameters of the Whisper Large V3 model using the CHiME-8 training data, with an early stopping set to a patience of 5,000 steps.
To obtain optimal performance of the CHiME-8 task, we fine-tuned the Whisper Large V3 model on the CHiME-8 training data. However, the CHiME-8 training data is very noisy and part of it is unreliable for fine-tuning. We adopted a curriculum learning scheme to mitigate this issue. We proposed to filter out utterances that have a high character error rate (CER) and can thus be considered unreliable. Here, we computed the CER on the fly during training. This means that we can gradually increase the number of training utterances as fine-tuning progresses and the model becomes stronger and CER decreases.
To implement such a curriculum learning practically, we changed the transcription reference to the self-generated decoding result when the CER exceeds 30 %, and multiplied the loss by a small value of $10^{-3}$ to reduce its impact.
In preliminary experiments, we found that the proposed scheme could stabilize training. However, we could not investigate its impact deeply due to the limited time required to develop the challenge system and the cost of fine-tuning the Whisper Large model. This model turned out to be our best ASR model, as shown in the ablation study of Table 10 in A.3.
@@ 5.2.2 Whisper Medium (ASR2)
The Whisper Medium English model has approximately 770M parameters and consists of a 24-layer Transformer encoder and decoder, each with 8 attention heads and a model width of 1024. The large model uses the byte-level BPE tokenizer of GPT-2 and has the same vocabulary size of 51,864 tokens.
We fined-tuned all model parameters using the CHiME-8 training data, with an early stopping set to a patience of 5,000 steps. Note that this model was created in the early development of our system without the curriculum learning scheme of ASR1. As shown in our ablation study of A.3, it performed worse than ASR1 but was still effective for system combination.
@@ 5.2.3 NeMo Transducer (ASR3)
The CHiME baseline employs a NeMo transducer model [74] for the ASR backend. This model has 644M parameters, i.e., two-layer 2D convolutional CNNs followed by 24 fast Conformer blocks [75, 76]. The prediction and joint networks consist of a 640-dimensional long short-term memory (LSTM) and a 640-dimensional feed-forward network. The vocabulary consists of 1025 BPE tokens.
@@ 5.2.4 WavLM Transducer (ASR4)
Although other ASR systems (ASR1-3) and the WavLM pre-encoder are pre-trained using a large amount of data, ASR4 contains randomly initialized parameters, i.e., the branchformer encoder, prediction, and joint networks, which are trained from scratch. These network parameters should also be trained on a large amount of data. For this model training, we used supervised data from CHiME-8 and LibriSpeech, as well as unsupervised data from the vast amount of VoxCeleb1+2 dataset [51]. We used VoxCeleb1+2, transcribed by a publicly available pre-trained NeMo CTC model, as semi-supervised training data. To discard noisy data from the large VoxCeleb1+2 dataset, we performed data selection for efficient training. Moreover, we also conducted multi-step training, which drastically improved the performance of the integration of the WavLM and transducer model. A more detailed procedure for data selection and multi-step training is described as follows:
@@
    Data selection The contrastive data selection algorithm [35, 19] is applied to the VoxCeleb1+2 data. This algorithm employs a general and a target LM to select data with high domain relevance scores on the target domain and low scores on the general domain. Each LM consists of a 5-gram LM with Kneser-Ney smoothing, trained using discrete acoustic units on the general (VoxCeleb1+2) and target (CHiME-8) domain datasets, respectively. In this work, to generate the acoustic units, we discretized the 21st WavLM layer output instead of using the w2v-BERT codebook [35]. For the quantization, we run the k-means algorithm with 1024 clusters. We calculated the domain relevance score by measuring the probability differences between the target- and general-domain LMs.
@@
    Multi-step training: In previous work [19], we found a novel training scheme that consisted of multiple training steps. We conducted three-step training to build the ASR4 system sequentially: 1) partial parameters were trained using the CHiME & LibriSpeech & VoxCeleb1+2 datasets while freezing WavLM frontend, 2) all network parameters (including WavLM) were finetuned using the same data from the first step, and 3) we finetuned it using only the CHiME-8 data. Note that we used early stopping with patience of 5 epochs for all training steps.
@@
The ASR4 model achieved competitive performance with the larger ASR1 model, as shown in our experiments of A.3.
@@ 5.3 LM rescoring and system combination (ASR5)
where $\hat{\mathbf{x}}_{s,j}$ is the input speech obtained from SE module, $\hat{\mathbf{v}}_{s,j}$ is the hypothesis, $\text{log}\ p_{\text{ASR}}(\mathbf{v}_{s,j}|\hat{\mathbf{x}}_{s,j})$ is the ASR model score for $\mathbf{v}_{s,j}$ given $\hat{\mathbf{x}}_{s,j}$, which corresponds to Eq. (5). $\text{log}\ p_{\text{LM}}(\mathbf{v}_{s,j})$ indicates the LM score for $\mathbf{v}_{s,j}$, $\beta_{1}$ is the weight for LM, $\beta_{2}$ is the reward given proportional to the length of $\mathbf{v}_{s,j}$ (the weight $\beta_{1}$ and reward $\beta_{2}$ were tuned using the development set), and $\hat{\mathbf{v}}_{s,j}$ is the best hypothesis.
We built a Transformer-LM consisting of 35M parameters for LM rescoring. The LM has the vocabulary of 1000 BPE tokens. We pre-trained the LM using 1/10 of the LibriSpeech text dataset and then finetuned using the CHiME-8 train text dataset. At the inference, we use 256 past rescored (re-ranked) 1-best tokens as the context to rescore the current hypotheses (i.e., context carry-over) [78, 79].
We built six different ASR models (the four models described above and two versions of ASR1 and ASR4 with different training setups). These systems vary in architecture, training procedure, and data. Therefore, they show different error patterns. To improve performance, we performed the system combination of the original 1-best hypotheses of the four ASR systems (ASR5) and their rescored versions by an external LM, using ROVER [32], as shown in Fig. 9. We show the effectiveness of system combination in the experiments of A.3.
@@ 6 Related Works
We briefly summarize STCON CHiME-8 system [20], which achieved top performance on the DASR task. STCON follows a diarization-first pipeline and also uses TS-VAD.
Diarization: For pre-diarization, they use VAD to find speech activity combined with a more advanced speaker embedding extractor, which provides as many embeddings as speakers in a chunk. These embeddings are combined with ECAPA-TDNN embeddings. For clustering, they used GMM-based clustering, which is also used for speaker counting. To further refine speaker counting, they developed a Wav2Vec2.0-based classifier to detect if speech from a cluster corresponds to non-speech and should thus be discarded. They perform microphone selection before diarization.
@@
ASR:: Several backends were developed, incorporating various types of ASR systems: 1) TDNN-based acoustic models within DNN-HMM hybrid ASR systems trained using Kaldi [80]; 2) three AED models, specifically Uconv-Conformer [81], E-Branchformer [77], and Whisper, all trained with ESPnet [82]; and 3) a Transducer model utilizing the Zipformer encoder [83], trained with K2999https://github.com/k2-fsa. Note that all ASR systems employed WavLM Large as input features. ASR was performed on two versions of the enhanced speech using these backends, and their results were combined. Besides, they also used session-wise unsupervised adaptation.
To summarize, our system’s major difference is its use of more advanced speaker embedding extraction, which may achieve better speaker counting. The SE part uses an additional neural TSE frontend. Finally, they combine many diarization systems and many ASR systems to obtain their final results.
@@ 7 Experiments
We report results on the CHiME-8 development and evaluation sets used during the challenge. The test set comprises four datasets: CHiME-6, DiPCO, Mixer 6, and NOTSOFAR-1. CHiME-6 and DIPCO consist of four-speaker home-party recordings of about 2.5 hours and 30 minutes, respectively. Mixer 6 data are 2-speaker interviews that are 15 minutes long, and NOSOFAR1 data are short 6-minute meetings of 4 to 8 participants recorded in office meeting rooms. We refer to the challenge overview paper for more details about the data [53].
@@ 7.1 Evaluated systems
We report results with different versions of our system, which we compare to the NeMo baseline and the STCON system [20].
For diarization, we tested a system using EEND-VC without TS-VAD refinement (DIA1), described in Section 3.1, and a system with TS-VAD refinement (DIA2), described in section 3.4, which were developed for the challenge. We also experimented with a third diarization system, which uses a TS-VAD model trained with more data (DIA3). Besides, we also explored the impact of speaker counting errors by comparing results using DIA3 with Oracle speaker counts in our diarization pipeline, which we refer to as DIA4.
For SE, except when explicitly mentioned, all our systems used the same SE module described in Section 4, with improved microphone subset selection and beamformer.
For ASR, we report results using a WavLM-based transducer (ASR4), see Section 5.2.4, the lightest yet competitive ASR model; Whisper-L (ASR1), described in Section 5.2.4, a larger and more powerful system; and the system combination of the output of six ASR systems (ASR1 and ASR2 plus two variants of ASR1 and ASR4 with different training hyperparameters) with and without LM rescoring (ASR5), amounting to a total of 12 systems, as described in Section 5.3.
@@ 7.2 Evaluation metrics
The systems submitted to the CHiME-8 challenge were evaluated in terms of tcpWER [38], which measures WER between the reference transcription of each speaker and the hypotheses of the system, including speaker and time allocation, but allowing a 5-second margin between the reference and hypotheses.
To further analyze the results, we also report diarization performance in terms of DER and speaker counting performance. For the latter, we report the speaker counting accuracy (SCA) and the mean absolute speaker counting error (SCE), which is the mean of the absolute difference between the estimated and actual number of speakers.
@@ 7.3 Overall results
Table 1: Overall results on the development set in terms of DER [%] ($\downarrow$) SCE ($\downarrow$) and tcpWER [%] ($\downarrow$). DER is computed with md-eval with a collar of 0.25 sec. The best performances among all systems are indicated in bold fonts. The best performances among our systems are underlined. —”—  denotes a ditto mark. All our systems use the SE frontend described in Section 4.
@@
Table 2: Overall results on the evaluation set in terms of DER [%] ($\downarrow$) SCE ($\downarrow$) and tcpWER [%] ($\downarrow$). DER is computed with md-eval with a collar of 0.25 sec. The best performances among all systems are indicated in bold fonts. The best performances among our systems are underlined. —”—  denotes a ditto mark. All our systems use the SE frontend described in Section 4.
@@
We first report overall results in terms of DER, SCE, and tcpWER computed on the different datasets and their macro average. The reader can find ablation studies on the different modules in A.
@@
Tables 1 and 2 show the results on the development and evaluation sets, respectively. We compare our results with the challenge’s NeMo baseline [3] (system 0) and STCON (unconstrained_lm) [^10] STCON (unconstrained_lm) system was submitted to the unconstrained LM track of the DASR task, which allows using large language models (LLMs). It uses Llama-2-7B7 LLM for LM restoring as in [79]. LM rescoring with LLM improved their performance by 0.3 points in terms of Macro tcpWER on the evaluation set. Our systems do not employ such LLM-based LM rescoring. system [20](system 0’) that achieved the best tcpWER on the task at the challenge. [^11] https://www.chimechallenge.org/challenges/chime8/task1/results We report the results of the systems we submitted to the challenge, as well as post-challenge results. We submitted three systems to the challenge, system 1 is the most lightweight model using only the simpler DIA1 and ASR4 models, system 2 uses our best diarization (DIA2) [^12] At the time of the challenge submission and ASR (ASR1) models, and system 3 combines the recognition hypotheses obtained with several ASR backends (ASR5).
@@
We also report post-challenge results (systems 4 and 5), where we replaced the diarization module (DIA2) of systems 2 and 3 with a stronger diarization system (DIA3), which uses a TS-VAD model trained with more data. The EEND model, speaker counting, SE, and ASR modules remain the same. Note that all our systems rely on the same speaker counting module, and consequently, the estimated number of speakers is the same.
@@
We also show results with the Oracle number of speakers (systems 6 and 7), where we replaced the estimated speaker counts in DIA3 of systems 4 and 5 with the Oracle ones (DIA4) (all other components remain the same). Indeed, we believe that for some practical applications such as meeting transcriptions, it is reasonable to assume that we can know in advance the total number of speakers. Besides, it helps us understand the impact of speaker counting errors on the overall performance.
@@
Among our systems (without Oracle information), naturally, the most advanced systems, 3 and 5, achieve the best macro DER and tcpWER for the challenge and post-challenge, respectively. Our systems greatly reduce the tcpWER over the challenge baseline for all test conditions and achieve performance close to the STCON system. Notably, our stronger post-challenge diarization system improved our challenge macro DER by more than three points, outperforming all challenge submissions. This resulted in more than one point reduction of the macro tcpWER results, also achieving the best performance on the development and evaluation sets of NOTSOFAR-1 set. Comparing the results on the development and evaluation sets, we observe that our system has a higher SCE on the evaluation set. This may indicate over-tuning the parameters on the development set.
@@
Figure 10: Impact of speaker count errors on tcpWER of the different sessions of the evaluation set. The vertical axis shows the difference between the tcpWER of systems using estimated and Oracle speaker counts, i.e., systems 5 (DIA3, SE, ASR5) and 7 (DIA4(oracle speaker nb) SE ASR5), respectively. We report results only for the sessions with speaker counting errors. The number above each bar indicates the speaker counting error. Positive values represent false alarms, while negative values represent missed speakers.
@@
Fig. 10 shows the impact of speaker counting errors on tcpWER by comparing results without and with Oracle speaker counts for all sessions of the evaluation set. We observe that for some sessions, wrongly estimating the number of speakers can have dramatic consequences in terms of tcpWER. Indeed, if a speaker is split into two speakers, it will lead to a large number of deletion and insertion errors. For example, for Mixer 6, we only overestimated the number of speakers for two of the 23 sessions in the eval set. However, for one, it led to a drastic increase of the tcpWER reaching about 85 % because one of the speakers was split into two speakers, but the speech was well transcribed. The error on this session alone explains the relatively high tcpWER we obtained on that dataset compared with when using the Oracle number of speakers.
@@
In contrast, for several sessions, speaker errors degrade tcpWER by less than 5 points and can sometimes even lead to slightly better performance, as seen in Fig. 10. This shows that SCEs do not always have a high impact on performance. This is probably the case when the amount of speech allocated to the wrong speaker is small.
@@
The results with the Oracle number of speakers show that further improving speaker counting accuracy is crucial to achieving better performance with our system. In addition, although the tcpWER level on tasks such as Mixer 6 and NOTSOFAR-1 are already relatively low, they remain above 20 % on tasks such as CHiME-6 and DipCo. This indicates that further progress in diarization, SE, and ASR would be needed to bring down the tcpWER in challenging home-party recording scenarios.

### v2

@@ 5 Speech recognition backend
Figure 9: Proposed ASR backend.
In this section, we describe how we implemented our ASR backend (“$\text{ASR}(\cdot)$” in Eq. (5)). Figure 9 shows the schematic diagram of our ASR backend, which uses four different ASR models, performs language model (LM) for restoring, and combines their output with system combination. The details of the ASR models, the training data used for their training, the language model (LM) used for rescoring, and the system combination approach are described below.
@@ 5.1 Training data for ASR backend
We used 70 hours of CHiME-8 training data released by the challenge organizers[56] (including data from CHiME-6, DiPCo, Mixer 6, and NOTSOFAR-1 (only the real recordings)) processed with GSS for the Oracle segmentation, which is commonly used for building ASR models (ASR1-4). We did not use train_call and train_intv in Mixer 6 at this time because preliminary experiments showed minimal improvements with them. For pre-training of the transducer-based ASR model with WavLM Large (ASR4), we also used LibriSpeech [58] and VoxCeleb1+2 [54] datasets in addition to the CHiME-8 described above. The development set was utilized for measuring token accuracy or training loss for early stopping.
@@ 5.2 ASR models
We implemented four ASR models: two attentional encoder-decoder (AED) models [73] and two transducer models [74]. For building strong ASR systems, we adopt several pre-trained models: 1) two different sizes of Whisper-based AED models (ASR1/ASR2), 2) the official pre-trained transducer model based on the CHiME-8 DASR NeMo baseline system (ASR3), and 3) the self-supervised learning (SSL) model, i.e., WavLM [52], for the encoder of the transducer model (ASR4). The following describes the detailed model structure and training pipeline of each ASR model.
@@ 5.2.1 Whisper Large v3 (ASR1)
Whisper Large v3 model [75] is a strong ASR model trained on a large amount of data. It consists of 1540M parameters, i.e., a 32-layer Transformer encoder and decoder, each with 8 attention heads and a hidden dimension of 1280. It utilizes the byte-level byte-pair encoding (BPE) [76] tokenizer from GPT-2 [77] and has a vocabulary size of 51,864 tokens. We fine-tuned all parameters of the Whisper Large v3 model using the CHiME-8 training data, with an early stopping set to a patience of 5,000 steps.
To obtain an optimal performance on the CHiME-8 task, we fine-tuned the Whisper Large v3 model on the CHiME-8 training data. However, since this data is very noisy and part of it is unreliable for fine-tuning, we also adopted a curriculum learning scheme. Our idea is to filter out utterances that have a high character error rate (CER) and can thus be considered unreliable. Here, we computed the CER on the fly during training. This means that we can gradually increase the number of training utterances as fine-tuning progresses, the model becomes stronger and CER decreases.
To implement this curriculum learning practically, we changed the transcription reference to the self-generated decoding result when the CER exceeds 30 %, and multiplied the loss by a small value of $10^{-3}$ to reduce its impact.
In preliminary experiments, we found that the proposed scheme could stabilize training. However, we could not investigate its impact deeply due to the limited time required to develop the challenge system and the cost of fine-tuning the Whisper Large model. This model turned out to be our best ASR model, as shown in the ablation study in A.3(Table 10).
@@ 5.2.2 Whisper Medium (ASR2)
The Whisper Medium English model has approximately 770M parameters and consists of a 24-layer Transformer encoder and decoder, each with 8 attention heads and a model width of 1024. The large model utilizes the byte-level BPE tokenizer of GPT-2 and has the same vocabulary size of 51,864 tokens.
We fined-tuned all model parameters using the CHiME-8 training data, with an early stopping set to a patience of 5,000 steps. Note that this model was created in the early development of our system without the curriculum learning scheme of ASR1. As shown in our ablation study (A.3), it performed worse than ASR1 but was still effective for system combination.
@@ 5.2.3 NeMo Transducer (ASR3)
The CHiME baseline utilizes a NeMo transducer model [78] for the ASR backend. This model has 644M parameters, i.e., two-layer 2D convolutional CNNs followed by 24 fast Conformer blocks [79, 80]. The prediction and joint networks consist of a 640-dimensional long short-term memory (LSTM) and a 640-dimensional feed-forward network. The vocabulary consists of 1025 BPE tokens.
@@ 5.2.4 WavLM Transducer (ASR4)
Although the other ASR systems (ASR1-3) and the WavLM pre-encoder are pre-trained using a large amount of data, ASR4 contains randomly initialized parameters, i.e., the branchformer encoder, prediction, and joint networks, which are trained from scratch. These network parameters should also be trained on a large amount of data. For this model training, we used supervised data from CHiME-8 and LibriSpeech, as well as unsupervised data from the vast amount of VoxCeleb1+2 dataset [54]. We utilized VoxCeleb1+2, transcribed by a publicly available pre-trained NeMo CTC model, as semi-supervised training data. To discard noisy data from the large VoxCeleb1+2 dataset, we performed data selection for efficient training. We also conducted multi-step training, which dramatically improved the performance of the integration of the WavLM and transducer model. We describe the data selection and multi-step training processes below.
@@
    Data selection The contrastive data selection algorithm [35, 19] is applied to the VoxCeleb1+2 data. This algorithm utilizes a general LM and a target LM to select data with high domain relevance scores on the target domain and low scores on the general domain. Each LM consists of a 5-gram LM with Kneser-Ney smoothing, trained using discrete acoustic units on the general (VoxCeleb1+2) and target (CHiME-8) domain datasets, respectively. In this work, to generate the acoustic units, we discretized the 21st WavLM layer output instead of using the w2v-BERT codebook [35]. For the quantization, we run the k-means algorithm with 1024 clusters. We calculated the domain relevance score by measuring the probability differences between the target- and general-domain LMs.
@@
    Multi-step training: In previous work [19], we found a novel training scheme that consists of multiple training steps. We conducted three-step training to build the ASR4 system sequentially: 1) partial parameters were trained using the CHiME & LibriSpeech & VoxCeleb1+2 datasets while freezing the WavLM frontend, 2) all network parameters (including WavLM) were fine-tuned using the same data from the first step, and 3) we fine-tuned it using only the CHiME-8 data. Note that we used early stopping with patience of 5 epochs for all training steps.
@@
The ASR4 model achieved a competitive performance with the larger ASR1 model, as shown in our experiments in A.3.
@@ 5.3 LM rescoring and system combination (ASR5)
where $\hat{\mathbf{x}}_{s,j}$ is the input speech obtained from the SE module, $\hat{\mathbf{v}}_{s,j}$ is the hypothesis, $\text{log}\ p_{\text{ASR}}(\mathbf{v}_{s,j}|\hat{\mathbf{x}}_{s,j})$ is the ASR model score for $\mathbf{v}_{s,j}$ given $\hat{\mathbf{x}}_{s,j}$, which corresponds to Eq. (5). $\text{log}\ p_{\text{LM}}(\mathbf{v}_{s,j})$ indicates the LM score for $\mathbf{v}_{s,j}$, $\beta_{1}$ is the weight for the LM, $\beta_{2}$ is the reward given proportional to the length of $\mathbf{v}_{s,j}$ (the weight $\beta_{1}$ and reward $\beta_{2}$ were tuned using the development set), and $\hat{\mathbf{v}}_{s,j}$ is the best hypothesis.
We built a Transformer-LM consisting of 35M parameters for LM rescoring. The LM has the vocabulary of 1000 BPE tokens. We pre-trained the LM using 1/10 of the LibriSpeech text dataset and then fine-tuned using the CHiME-8 train text dataset. At the inference, we used 256 past rescored (re-ranked) 1-best tokens as the context to rescore the current hypotheses (i.e., context carry-over) [82, 83].
We built six different ASR models (the four models described above and two versions of ASR1 and ASR4 with different training setups). These systems vary in architecture, training procedure, and data. Therefore, they show different error patterns. To improve performance, we performed the system combination of the original 1-best hypotheses of the four ASR systems (ASR5) and their rescored versions by an external LM, using ROVER [32], as shown in Figure 9. We discuss the effectiveness of the system combination in the experiments in A.3.
@@ 6 Related works
We briefly summarize the STCON CHiME-8 system [20], which achieved the top performance on the DASR task. STCON follows a diarization-first pipeline and also uses TS-VAD.
Diarization: For pre-diarization, they use VAD to find speech activity combined with a more advanced speaker embedding extractor, which provides as many embeddings as speakers in a chunk. These embeddings are combined with ECAPA-TDNN embeddings. For clustering, they used GMM-based clustering, which is also utilized for speaker counting. To further refine speaker counting, they developed a Wav2Vec2.0-based classifier to detect if speech from a cluster corresponds to non-speech and should thus be discarded. Microphone selection is performed before diarization.
@@
ASR:: Several backends has been developed, including 1) TDNN-based acoustic models within DNN-HMM hybrid ASR systems trained using Kaldi [84]; 2) three AED models, specifically Uconv-Conformer [85], E-Branchformer [81], and Whisper, all trained with ESPnet [86]; and 3) a Transducer model utilizing the Zipformer encoder [87], trained with K2999https://github.com/k2-fsa. All of these ASR systems utilized WavLM Large as input features. ASR was performed on two versions of the enhanced speech using these backends, and their results were combined. They also used session-wise unsupervised adaptation.
To summarize, the major difference of our system compared to the above is its use of more advanced speaker embedding extraction, which may achieve better speaker counting. The SE part uses an additional neural TSE frontend. Finally, they combine many diarization systems and many ASR systems to obtain their final results.
@@ 7 Experiments
In this section, we report the results on the CHiME-8 development and evaluation sets used during the challenge. The test set comprises four datasets: CHiME-6, DiPCo, Mixer 6, and NOTSOFAR-1. CHiME-6 and DiPCo consist of four-speaker home-party recordings of about 2.5 hours and 30 minutes, respectively. Mixer 6 data are 2-speaker interviews that are 15 minutes long, and NOTSOFAR-1 data are short 6-minute meetings of 4 to 8 participants recorded in office meeting rooms. We refer to the challenge overview paper for more details about the data [56].
@@ 7.1 Evaluated systems
We report the results with different versions of our system, which we compare to the NeMo baseline and the STCON system [20].
For diarization, we tested a system using EEND-VC without TS-VAD refinement (DIA1), described in Section 3.1, and a system with TS-VAD refinement (DIA2), described in section 3.4, which were developed for the challenge. We also experimented with a third diarization system, that uses a TS-VAD model trained with more data (DIA3). In addition, we also explored the impact of speaker counting errors by comparing the results using DIA3 with Oracle speaker counts in our diarization pipeline, which we refer to as DIA4.
For SE, except when explicitly mentioned, all our systems utilized the same SE module described in Section 4, with improved microphone subset selection and beamformer.
For ASR, we report the results using a WavLM-based transducer (ASR4) (see Section 5.2.4), the lightest yet competitive ASR model; Whisper-L (ASR1), described in Section 5.2.4, a larger and more powerful system; and the system combination of the output of six ASR systems (ASR1 and ASR2 plus two variants of ASR1 and ASR4 with different training hyperparameters) with and without LM rescoring (ASR5), amounting to a total of 12 systems, as described in Section 5.3.
@@ 7.2 Evaluation metrics
The systems submitted to the CHiME-8 challenge were evaluated in terms of tcpWER [38], which measures the WER between the reference transcription of each speaker and the hypotheses of the system, including speaker and time allocation, but allowing a 5-second margin between the reference and hypotheses.
To further analyze the results, we also report the diarization performance in terms of DER and speaker counting performance. For the latter, we report the speaker counting accuracy (SCA) and the mean absolute speaker counting error (SCE), which is the mean of the absolute difference between the estimated and actual number of speakers.
@@ 7.3 Overall results
Table 1: Overall results on the development set in terms of DER [%] ($\downarrow$) SCE ($\downarrow$) and tcpWER [%] ($\downarrow$). DER is computed using md-eval with a collar of 0.25 sec. The best performances among all systems are indicated in bold fonts, and the best performances among our systems are underlined. —”—  denotes a ditto mark. All our systems use the SE frontend described in Section 4.
@@
Table 2: Overall results on the evaluation set in terms of DER [%] ($\downarrow$) SCE ($\downarrow$) and tcpWER [%] ($\downarrow$). DER is computed using md-eval with a collar of 0.25 sec. The best performances among all systems are indicated in bold fonts, and the best performances among our systems are underlined. —”—  denotes a ditto mark. All our systems use the SE frontend described in Section 4.
@@
We first report the overall results in terms of DER, SCE, and tcpWER computed on the different datasets and their macro average. Ablation studies on the different modules are presented in A.
@@
Tables 1 and 2 show the results on the development and evaluation sets, respectively. We compare our results with the challenge’s NeMo baseline [3] (system 0) and STCON (unconstrained_lm) [^10] STCON (unconstrained_lm) system was submitted to the unconstrained LM track of the DASR task, which allows the use of large language models (LLMs). It utilizes a Llama-2-7B7 LLM for LM restoring as in [83]. LM rescoring with LLM improved the performance by 0.3 points in terms of Macro tcpWER on the evaluation set. Our systems do not utilize such LLM-based LM rescoring. system [20](system 0’) that achieved the best tcpWER on the task at the challenge. [^11] https://www.chimechallenge.org/challenges/chime8/task1/results We report the results of the three systems we submitted to the challenge: system 1, which is the most lightweight model and uses only the simpler DIA1 and ASR4 models, system 2, which utilizes our best diarization (DIA2) [^12] At the time of the challenge submission. and ASR (ASR1) models, and system 3, which combines the recognition hypotheses obtained with several ASR backends (ASR5).
@@
We also report the post-challenge results (systems 4 and 5), where we replaced the diarization module (DIA2) of systems 2 and 3 with a stronger diarization system (DIA3), which uses a TS-VAD model trained with more data. The EEND model, speaker counting, SE, and ASR modules remain the same. Note that all our systems rely on the same speaker counting module, and consequently, the estimated number of speakers is the same.
@@
We also show results with the Oracle number of speakers (systems 6 and 7), where we replaced the estimated speaker counts in DIA3 of systems 4 and 5 with the Oracle ones (DIA4) (all other components remain the same). We believe that for some practical applications such as meeting transcriptions, it is reasonable to assume that we can know in advance the total number of speakers. Besides, it helps us understand the impact of speaker counting errors on the overall performance.
@@
Among our systems (without Oracle information), the most advanced systems, 3 and 5, naturally achieve the best macro DER and tcpWER for the challenge and post-challenge, respectively. Our systems greatly reduce the tcpWER over the challenge baseline for all test conditions and achieve performance close to the STCON system. Notably, our stronger post-challenge diarization system improved our challenge macro DER by more than three points, outperforming all challenge submissions. This resulted in a reduction of more than one point reduction of the macro tcpWER results, also achieving the best performance on the development and evaluation sets of NOTSOFAR-1 set. Comparing the results on the development and evaluation sets, we observe that our system has a higher SCE on the evaluation set. This may indicate over-tuning of the parameters on the development set.
@@
Figure 10: Impact of speaker count errors on tcpWER of the different sessions of the evaluation set. Vertical axis shows the difference between the tcpWER of systems using estimated and Oracle speaker counts, i.e., systems 5 (DIA3, SE, ASR5) and 7 (DIA4(Oracle speaker nb) SE ASR5), respectively. We report results only for sessions with speaker counting errors. The number above each bar indicates the speaker counting error. Positive values represent false alarms, while negative values represent missed speakers.
@@
Figure 10 shows the impact of speaker counting errors on tcpWER by comparing the results without and with Oracle speaker counts for all sessions of the evaluation set. We observe that for some sessions, wrongly estimating the number of speakers can have dramatic consequences in terms of tcpWER. Indeed, if a speaker is split into two speakers, it will lead to a large number of deletion and insertion errors. For example, for Mixer 6, we only overestimated the number of speakers for two of the 23 sessions in the eval set. However, for one, it led to a drastic increase of the tcpWER reaching about 85 % because one of the speakers was split into two speakers, but the speech was well transcribed. The error on this session alone explains the relatively high tcpWER we obtained on that dataset compared to when using the Oracle number of speakers.
@@
In contrast, for several sessions, speaker errors degrade the tcpWER by less than 5 points and can sometimes even lead to slightly better performance, as seen in Figure 10. This indicates that SCEs do not always have a high impact on performance. This is probably the case when the amount of speech allocated to the wrong speaker is small.
@@
The results with the Oracle number of speakers show that further improvement to the speaker counting accuracy is crucial to achieving better performance with our system. In addition, although the tcpWER level on tasks such as Mixer 6 and NOTSOFAR-1 are already relatively low, they remain above 20 % on tasks such as CHiME-6 and DiPCo. This indicates that further progress in diarization, SE, and ASR would be needed to bring down the tcpWER in challenging home-party recording scenarios.